# Video Understanding with Large Language Models: A Survey

### v4

@@ Abstract
With the burgeoning growth of online video platforms and the escalating volume of video content, the demand for proficient video understanding tools has intensified markedly. Given the remarkable capabilities of large language models (LLMs) in language and multimodal tasks, this survey provides a detailed overview of recent advancements in video understanding that harness the power of LLMs (Vid-LLMs). The emergent capabilities of Vid-LLMs are surprisingly advanced, particularly their ability for open-ended multi-granularity (abstract, temporal, and spatiotemporal) reasoning combined with commonsense knowledge, suggesting a promising path for future video understanding. We examine the unique characteristics and capabilities of Vid-LLMs, categorizing the approaches into three main types: Video Analyzer $\times$ LLM, Video Embedder $\times$ LLM, and (Analyzer + Embedder) $\times$ LLM. Furthermore, we identify five sub-types based on the functions of LLMs in Vid-LLMs: LLM as Summarizer, LLM as Manager, LLM as Text Decoder, LLM as Regressor, and LLM as Hidden Layer. Furthermore, this survey presents a comprehensive study of the tasks, datasets, benchmarks, and evaluation methodologies for Vid-LLMs. Additionally, it explores the expansive applications of Vid-LLMs across various domains, highlighting their remarkable scalability and versatility in real-world video understanding challenges. Finally, it summarizes the limitations of existing Vid-LLMs and outlines directions for future research. For more information, readers are recommended to visit the repository at https://github.com/yunlong10/Awesome-LLMs-for-Video-Understanding.
@@ I Introduction
We live in a multimodal world where video has become the predominant form of media. With the rapid expansion of online video platforms and the growing prevalence of cameras in surveillance, entertainment, and autonomous driving, video content has risen to prominence as a highly engaging and rich medium, outshining traditional text and image-text combinations in both depth and appeal. This advancement has promoted an exponential surge in video production, with millions of videos being created daily. However, manually processing such a sheer volume of video content is labor-intensive and time-consuming. Consequently, there is a growing demand for tools to effectively manage, analyze, and process this abundance of video content. To meet that, video understanding methods and analysis technologies have emerged, leveraging intelligent analysis techniques to automatically recognize and interpret video content, thus significantly reducing the workload on human operators. Moreover, the continuous advancements in these methods enhance their task-solving capabilities, enabling them to handle a diverse range of video understanding tasks with increasing proficiency.
@@ I-A Development of Video Understanding Methods
As shown in Figure 1, the evolution of video understanding methods can be divided into four stages:
@@ I-A1 Conventional Methods
In the early stages of video understanding, handcrafted feature extraction techniques such as Scale-Invariant Feature Transform (SIFT) [1], Speeded-Up Robust Features (SURF) [2], and Histogram of Oriented Gradients (HOG) [3] were used to capture key information in videos. Background Subtraction [4], optical flow methods [5], and Improved Dense Trajectories (IDT) [6, 7] were used to model the motion information for tracking. Since videos can be viewed as time-series data, temporal analysis techniques like Hidden Markov Models (HMM) [8] were also used to understand video content. Before the popularity of deep learning, basic machine learning algorithms such as Support Vector Machines (SVM) [9], Decision Trees [10], and Random Forests were also used in video classification and recognition tasks. Cluster analysis [11] for classifying video segments, or Principal Component Analysis (PCA) [12, 13] for data dimensionality reduction were also commonly used methods for video analysis.
@@ I-B Related Surveys
Previous survey papers either study specific sub-tasks in the area of video understanding or focus on methodologies beyond video understanding. For example, [78] surveys multimodal foundation models for general vision-language tasks, which includes both image and video applications. [79] and [80] focus on surveying video captioning and action recognition tasks, respectively. Other video understanding tasks, such as the video question answering and grounding, are not considered. Moreover, [81] and [82] survey video-related methodologies – video diffusion models and LLMs, respectively, lacking the concentration on video understanding. Despite the significant value to the community, previous survey papers leave a gap in surveying the general video understanding task based on large language models. This paper fills this gap by comprehensively surveying the video understanding task using large language models.
@@ I-C Survey Structure
This survey is structured as follows: Section II offers preliminaries for video understanding with LLMs, including a summary of various video understanding tasks that require handling different levels of granularity, their associated datasets, and evaluation metrics. The background of LLMs is also introduced in this section. In Section III, we delve into details of recent research leveraging LLMs for video understanding, presenting their unique approaches and impact in the field, where we divide these Vid-LLMs into three main categories, Video Analyzer $\times$ LLM and Video Embedder $\times$ LLM, and (Analyzer + Embedder) $\times$ LLM; and five sub-categories, LLM as Summarizer/Manager/Text Decoder/Regressor/Hidden Layer, shown as Figure 5. This section also includes the training strategies of Vid-LLMs. Section IV adds more information about popular ways to evaluate Vid-LLMs, together with some benchmarks and performances of some Vid-LLMs on the most commonly used benchmarks. Section V explores the application of Vid-LLMs across multiple significant fields. The survey concludes in Section VI by summarizing key findings and identifying unresolved challenges and potential areas for future research.
@@ II-A2 Temporal Understanding Tasks
  • Video Summarization: Video Summarization aims at condensing a long video into a shorter version while preserving essential content and important information. F1-score, Spearman, and Kendall usually evaluate this task as metrics. Commonly-used datasets include SumMe [112], TVSum [113], Ads-1k [114], VideoXum [98], Instrcut-V2Xum [100].
@@
  • Temporal Action/Event Localization: Temporal action/event localization task aims at identifying the precise temporal segments of actions or events within a video. By analyzing sequential frames, models trained for this task must indicate when specific activities start and end. Datasets for Temporal Action/Event Localization involve THUMOS’14 [117], ActivityNet-1.3 [84], and UnAV-100 [118].
@@
  • Generic Event Boundary Detection: Generic event boundary detection involves identifying certain frames in a video where significant changes or transitions occur and splitting videos based on different events or activities. Kinetics-GEBD [121] is a widely-used dataset for this task.
@@
  • Dense Video Captioning: Dense video captioning aims at generating detailed and continuous textual descriptions for multiple events and actions occurring throughout a video. Evaluation metrics like BLEU, METEOR, CIDEr, and ROUGE-L are used to evaluate this task. Relevant datasets are ActivityNet Captions [122], VidChapters-7M [123], YouCook2 [95], and ViTT [97].
@@ II-A3 Spatiotemporal Understanding Tasks
  • Object Tracking: Object tracking aims at continuously identifying and following the position of specific objects within a video over time. A good object tracking model should maintain accurate and consistent tracking trajectories of objects, even for videos with occlusions, changes in appearance, and motions. Benchmarks like OTB [124], UAV [125], and VOT [126] are commonly used for this task.
@@
  • Video Saliency Detection: Video saliency detection aims at identifying the most visually important and attention-grabbing regions in a video. This task highlights areas that stand out due to factors like motion, contrast, and unique features. Relevant datasets to this task are DHF1K [131], Hollywood-[85], UCF-Sports [132], AVAD [133], Coutrot1 [134], Coutrot2 [135], ETMD [136], and SumMe [112].
@@
  • Video Object Referring Segmentation: Video object referring segmentation involves segmenting specific objects in a video based on natural language descriptions. It identifies and isolates the referred objects accurately across frames, where MeViS [141] is a common benchmark for this task.
@@ II-B Background for LLMs
LLMs possess extensive generalization abilities that can be applied to various downstream tasks, including multi-modal tasks. Multimodal Large Language Models (MLLMs) [146, 76, 147, 148, 149] typically incorporate multimodal encoders, cross-modality aligners, and an LLM core structure. By combining multimodal encoders with the LLM core, MLLMs excel at integrating visual and linguistic contexts to produce detailed content.
@@ III Vid-LLMs
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=6em, edge+=darkgray, line width=1pt, s sep=20pt, inner xsep=3pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10.2em,font=,, where level=2text width=8.6em,font=,, where level=3text width=8.2em,font=,, where level=4text width=8.2em,font=,, [ Vid-LLMs, ver [ Video Analyzer $\times$ LLM [ LLM as Summarizer [ GIT-LLaVA [150], MM-Screenplayer [151], MoReVQA [152],
@@
VLog [161], LaViLa [162] , leaf, text width=27em ] ] [ LLM as Manager [ DrVideo [163], OmAgent [164], LVNet [165], GPTSee [166],
@@
Video ChatCaptioner [183], ChatVideo [183], ViperGPT [184] , leaf, text width=27em ] ] ] [ Video Embedder $\times$ LLM [ LLM as Text
Decoder [ Artemis [185], EmoLLM [186], FTFV-LLM [187],
@@
mPLUG-video [96], ChatBridge [240], Otter [241] , leaf, text width=28em ] ] [ LLM as Regressor [ Holmes-VAD [242], VideoLLM-online [243], VLM4HOI [244],
@@
GroundingGPT [251], TimeChat [252], VTimeLLM [253] , leaf, text width=27em ] ] [ LLM as Hidden
Layer [ VTG-LLM [254], VITRON [255], VTG-GPT [256],
@@
GPT4Video [260] , leaf, text width=23em ] ] ] [ (Analyzer + Embedder)
$\times$ LLM [ LLM as Manager [ MM-VID [261] , leaf, text width=7em ] ] [ LLM as Summarizer [ SUM-shot [262] , leaf, text width=7em ] ] [ LLM as Regressor [ Vriptor [263], Merlin [264], VideoChat [265], Vid2Seq [266] , leaf, text width=26.5em ] ] [ LLM as Text
Decoder [ Uni-AD [267], MM-narrator [268], Vamos [269], Auto-AD II [270] , leaf, text width=29em ] ] [ LLM as Hidden
Layer [ PG-Video-LLaVA [271] , leaf, text width=10em ] ] ] ]
@@ III-A1 Video Analyzer $\times$ LLM
  • LLM as Summarizer: In this subcategory, the primary function of the LLM is to summarize the analysis results obtained from the Video Analyzers. The summarization approach varies based on the prompts provided to the LLM, ranging from highly condensed summary texts and captions to comprehensive summaries for answering specific questions. Notably, in Video Analyzer $\times$ LLM as Summarizer systems, the information flow is usually unidirectional (see Figure 4), with data flowing from the video to the Video Analyzer and then to the LLM, without any reverse process. Examples of Vid-LLMs in the Video Analyzer $\times$ LLM as Summarizer category include: LaViLa [162], VLog [161], VAST [105], AntGPT [160], VIDOSC [159], Grounding-Prompter [158], LLoVi [157], Video ReCap [156], MVU [155], LangRepo [154], IG-VLM [153], MoReVQA [152], MM-Screenplayer [151], GIT-LLaVA [150], etc.
@@ III-A2 Video Embedder $\times$ LLM
A Video Embedder typically refers to a visual backbone/video encoder, such as ViT or CLIP, used to convert input videos into vector representations, known as video embeddings or video tokens. Some Embedders encode other modalities within the video, such as audio (e.g., CLAP), which are also categorized under Video Embedder here (note that we do not consider the LLM’s tokenizer as an embedder). Unlike the text generated by the Video Analyzer, the vectors generated by the Video Embedder cannot be directly utilized by the LLM and usually require an adapter to map these embeddings from the visual semantic space to the text semantic space of the LLM’s input tokens.
@@
This category of Vid-LLMs is relatively rare. As the name suggests, it involves simultaneously using a Video Analyzer to obtain textual analysis of the video and a Video Embedder to encode the video into embeddings. The LLM receives both types of inputs along with other prompts/instructions and outputs responses to complete tasks. The subcategories here can flexibly be any of the Summarizer/Manager/Text Decoder/Regressor/Hidden Layer categories. Examples of Vid-LLMs in the (Analyzer + Embedder) $\times$ LLM category include: Vid2Seq [266], VideoChat [265], MM-VID [261], Auto-AD II [270], Vamos [269], PG-Video-LLaVA [271], MM-Narrator [268], SUM-shot [262], Merlin [264], Uni-AD [267], Vriptor [263], etc.

### v5

@@ Abstract
With the rapid growth of online video platforms and the escalating volume of video content, the need for proficient video understanding tools has increased significantly. Given the remarkable capabilities of large language models (LLMs) in language and multimodal tasks, this survey provides a detailed overview of recent advances in video understanding that harness the power of LLMs (Vid-LLMs). The emergent capabilities of Vid-LLMs are surprisingly advanced, particularly their ability for open-ended multi-granularity (abstract, temporal, and spatiotemporal) reasoning combined with common-sense knowledge, suggesting a promising path for future video understanding. We examine the unique characteristics and capabilities of Vid-LLMs, categorizing the approaches into three main types: Video Analyzer $\times$ LLM, Video Embedder $\times$ LLM, and (Analyzer + Embedder) $\times$ LLM. We identify five subtypes based on the functions of LLMs in Vid-LLMs: LLM as Summarizer, LLM as Manager, LLM as Text Decoder, LLM as Regressor, and LLM as Hidden Layer. This survey also presents a comprehensive study of the tasks, datasets, benchmarks, and evaluation methods for Vid-LLMs. Additionally, it explores the extensive applications of Vid-LLMs in various domains, highlighting their remarkable scalability and versatility in real-world video understanding challenges. Additionally, it summarizes the limitations of existing Vid-LLMs and outlines directions for future research. For more information, readers are encouraged to visit the repository at https://github.com/yunlong10/Awesome-LLMs-for-Video-Understanding.
@@ I Introduction
We live in a multimodal world where video has become the predominant form of media. With the rapid expansion of online video platforms and the growing prevalence of cameras in surveillance, entertainment, and autonomous driving, video content has risen to prominence as a highly engaging and rich medium, outshining traditional text and image-text combinations in both depth and appeal. This advancement has fueled an exponential increase in video production, with millions of videos being created every day. However, manually processing such a sheer volume of video content is labor-intensive and time-consuming. As a result, there is a growing need for tools to effectively manage, analyze, and process this abundance of video content. To meet this need, video understanding methods have emerged that use intelligent analysis techniques to automatically recognize and interpret video content, significantly reducing the workload on human operators. In addition, the ongoing development of these methods is improving their task-solving capabilities, enabling them to handle a wide range of video understanding tasks with increasing proficiency.
@@ I-A Development of Video Understanding Methods
The evolution of video understanding methods can be divided into four stages, as shown in Figure 1:
@@ I-A1 Conventional Methods
In the early stages of video understanding, handcrafted feature extraction techniques such as Scale-Invariant Feature Transform (SIFT) [1], Speeded-Up Robust Features (SURF) [2], and Histogram of Oriented Gradients (HOG) [3] were used to capture key information in videos. Background Subtraction [4], optical flow methods [5], and Improved Dense Trajectories (IDT) [6, 7] were used to model the motion information for tracking. Since videos can be viewed as time series data, temporal analysis techniques such as Hidden Markov Models (HMM) [8] have also been used to understand video content. Before the popularity of deep learning, basic machine learning algorithms such as Support Vector Machines (SVM) [9], Decision Trees [10], and Random Forests were also used in video classification and recognition tasks. Cluster analysis [11] for classifying video segments, or Principal Component Analysis (PCA) [12, 13] for data dimensionality reduction have also been commonly used methods for video analysis.
@@ I-B Related Surveys
Previous survey papers either study specific sub-tasks in the area of video understanding or focus on methodologies beyond video understanding. For example, [78] surveys multimodal foundation models for general vision-language tasks, which includes both image and video applications. [79] and [80] focus on surveying video captioning and action recognition tasks, respectively. Other video understanding tasks, such as the video question answering and grounding, are not considered. Moreover, [81], [82], and [77] survey video-related methodologies, such as video diffusion models and LLMs, lacking the concentration on video understanding. [83] centers primarily on video foundation models, with insufficient attention given to language model-based approaches. Despite the significant value to the community, previous survey papers leave a gap in surveying the general video understanding task based on large language models. This paper fills this gap by comprehensively surveying the video understanding task using large language models.
@@ I-C Survey Structure
This survey is structured as follows: Section II offers preliminaries for video understanding with LLMs, including a summary of various video understanding tasks that require handling different levels of granularity, their associated datasets, and evaluation metrics. The background of LLMs is also introduced in this section. In Section III, we delve into details of recent research leveraging LLMs for video understanding, presenting their unique approaches and impact in the field, where we divide these Vid-LLMs into three main categories, Video Analyzer $\times$ LLM and Video Embedder $\times$ LLM, and (Analyzer + Embedder) $\times$ LLM; and five sub-categories, LLM as Summarizer/Manager/Text Decoder/Regressor/Hidden Layer, shown as Figure 5. This section also includes the training strategies of Vid-LLMs. Section IV adds more information about popular ways to evaluate Vid-LLMs, together with some benchmarks and performances of some Vid-LLMs on the most commonly used benchmarks. Section V explores the application of Vid-LLMs across multiple significant fields and identifies unresolved challenges and potential areas for future research.
@@ II-A2 Temporal Understanding Tasks
  • Video Summarization: Video Summarization aims at condensing a long video into a shorter version while preserving essential content. F1-score, Spearman, and Kendall usually evaluate this task as metrics. Commonly-used datasets include SumMe [112], TVSum [113], Ads-1k [114], VideoXum [98], Instrcut-V2Xum [100].
@@
  • Temporal Action/Event Localization: This task aims at identifying the precise temporal segments of actions or events within a video. By analyzing sequential frames, models trained for this task must indicate when specific activities start and end. Datasets for Temporal Action/Event Localization involve THUMOS’14 [117], ActivityNet-1.3 [85], and UnAV-100 [118].
@@
  • Generic Event Boundary Detection: Generic event boundary detection involves identifying certain frames in a video where significant changes occur and splitting videos based on different events or activities. Kinetics-GEBD [121] is a widely-used dataset for this task.
@@
  • Dense Video Captioning: Dense video captioning [122, 123, 124, 125, 126] aims at generating detailed and continuous textual descriptions for multiple events and actions occurring throughout a video. Evaluation metrics like BLEU, METEOR, CIDEr, and ROUGE-L are used to evaluate this task. Relevant datasets are ActivityNet Captions [127], VidChapters-7M [128], YouCook2 [95], and ViTT [97].
@@ II-A3 Spatiotemporal Understanding Tasks
  • Object Tracking: Object tracking aims at continuously identifying and following the position of specific objects within a video over time. A good tracking model should maintain accurate and consistent trajectories of objects, even for videos with occlusions, changes in appearance, and motions. Benchmarks like OTB [129], UAV [130], and VOT [131] are commonly used for this task.
@@
  • Video Saliency Detection: Video saliency detection aims at identifying the most visually important and attention-grabbing regions in a video [136]. This task highlights areas that stand out due to factors like motion, contrast, and unique features. Relevant datasets to this task are DHF1K [137], Hollywood-[86], UCF-Sports [138], AVAD [139], Coutrot1 [140], Coutrot2 [141], ETMD [142], and SumMe [112].
@@
  • Video Object Referring Segmentation: Video object referring segmentation involves segmenting specific objects in a video based on language descriptions. It identifies and isolates the referred objects accurately across frames, where MeViS [147] is a common benchmark for this task.
@@ II-B Background for LLMs
LLMs possess extensive generalization abilities that can be applied to various downstream tasks, including multi-modal tasks. Multimodal Large Language Models (MLLMs) [152, 76, 153, 154, 155] typically incorporate multimodal encoders, cross-modality aligners, and an LLM core structure. By combining multimodal encoders with the LLM, MLLMs excel at integrating visual and linguistic contexts to produce detailed content.
@@ III Vid-LLMs
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=6em, edge+=darkgray, line width=1pt, s sep=20pt, inner xsep=3pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10.2em,font=,, where level=2text width=8.6em,font=,, where level=3text width=8.2em,font=,, where level=4text width=8.2em,font=,, [ Vid-LLMs, ver, my-box, text width=12em [ Video Analyzer $\times$ LLM, analyzer-box, edge=myorange [ LLM as Summarizer, analyzer-box, edge=myorange [ GIT-LLaVA [156], MM-Screenplayer [157], MoReVQA [158],
@@
VLog [167], LaViLa [168] , leaf, text width=27em, edge=myorange, draw=myorange ] ] [ LLM as Manager, analyzer-box, edge=myorange [ DrVideo [169], OmAgent [170], LVNet [171], GPTSee [172],
@@
Video ChatCaptioner [189], ChatVideo [189], ViperGPT [190] , leaf, text width=27em, edge=myorange, draw=myorange ] ] ] [ Video Embedder $\times$ LLM, embedder-box, edge=mypink [ LLM as Text
Decoder, embedder-box, edge=mypink [ Artemis [191], EmoLLM [192], FTFV-LLM [193],
@@
mPLUG-video [96], ChatBridge [246], Otter [247] , leaf, text width=28em, edge=mypink, draw=mypink ] ] [ LLM as Regressor, embedder-box, edge=mypink [ Holmes-VAD [248], VideoLLM-online [249], VLM4HOI [250],
@@
GroundingGPT [256], TimeChat [257], VTimeLLM [258] , leaf, text width=27em, edge=mypink, draw=mypink ] ] [ LLM as Hidden
Layer, embedder-box, edge=mypink [ VTG-LLM [259], VITRON [260], VTG-GPT [261],
@@
GPT4Video [265] , leaf, text width=23em, edge=mypink, draw=mypink ] ] ] [ (Analyzer + Embedder)
$\times$ LLM, plus-box, edge=mygreen [ LLM as Manager, plus-box, edge=mygreen [ MM-VID [266] , leaf, text width=7em, draw=mygreen, edge=mygreen ] ] [ LLM as Summarizer, plus-box, edge=mygreen [ SUM-shot [267] , leaf, text width=7em, draw=mygreen, edge=mygreen ] ] [ LLM as Regressor, plus-box, edge=mygreen [ Vriptor [268], Merlin [269], VideoChat [270], Vid2Seq [271] , leaf, text width=26.5em, draw=mygreen, edge=mygreen ] ] [ LLM as Text
Decoder, plus-box, edge=mygreen [ Uni-AD [272], MM-narrator [273], Vamos [274], Auto-AD II [275] , leaf, text width=29em, draw=mygreen, edge=mygreen ] ] [ LLM as Hidden
Layer, plus-box, edge=mygreen [ PG-Video-LLaVA [276] , leaf, text width=10em, draw=mygreen, edge=mygreen ] ] ] ]
@@ III-A1 Video Analyzer $\times$ LLM
  • LLM as Summarizer: In this subcategory, the primary function of the LLM is to summarize the analysis obtained from the Video Analyzers. The summarization approach varies based on the prompts provided to the LLM, ranging from highly condensed summary texts and captions to comprehensive summaries for answering specific questions. Notably, in Video Analyzer $\times$ LLM as Summarizer systems, the information flow is usually unidirectional (see Figure 4), with data flowing from the video to the Video Analyzer and then to the LLM, without any reverse process. Examples of Vid-LLMs in the Video Analyzer $\times$ LLM as Summarizer category include: LaViLa [168], VLog [167], VAST [105], AntGPT [166], VIDOSC [165], Grounding-Prompter [164], LLoVi [163], Video ReCap [162], MVU [161], LangRepo [160], IG-VLM [159], MoReVQA [158], MM-Screenplayer [157], GIT-LLaVA [156], etc.
@@ III-A2 Video Embedder $\times$ LLM
A Video Embedder typically refers to a visual backbone/video encoder, such as ViT or CLIP, used to convert input videos into vector representations, known as video embeddings or video tokens. Some Embedders encode other modalities within the video, such as audio (e.g., CLAP [278]), which are also categorized under Video Embedder here (note that we do not consider the LLM’s tokenizer as an embedder). Unlike the text generated by the Video Analyzer, the vectors generated by the Video Embedder cannot be directly utilized by the LLM and usually require an adapter to map these embeddings from the vision’s (or other modalities’) semantic space to the text semantic space of the LLM’s input tokens.
@@
This category of Vid-LLMs is relatively rare. As the name suggests, it involves simultaneously using a Video Analyzer to obtain textual analysis of the video and a Video Embedder to encode the video into embeddings. The LLM receives both types of inputs along with other prompts/instructions and outputs responses to complete tasks. The subcategories here can flexibly be any of the Summarizer/Manager/Text Decoder/Regressor/Hidden Layer categories. Vid-LLMs in the (Analyzer + Embedder) $\times$ LLM category include: Vid2Seq [271], VideoChat [270], MM-VID [266], Auto-AD II [275], Vamos [274], PG-Video-LLaVA [276], MM-Narrator [273], SUM-shot [267], Merlin [269], Uni-AD [272], Vriptor [268], etc.